OpenStudy (caseysorensen):

Santino bought a book for $23.54. The price of the book was $22. What was the sales-tax rate? a.0.07 b.1.54 c.7 d.13

OpenStudy (caseysorensen):

@imqwerty

OpenStudy (caseysorensen):

i have no cule on this qusetion

imqwerty (imqwerty):

okay well the total price that he paid for the book must be the original price of the book + tax on the original price of book suppose the sales tax rate is \(x\)% then- total price paid= original price+ tax on original price \(23.54=22+ 22\times \large \frac{x}{100}\) try solving for \(x\) from here =]

OpenStudy (caseysorensen):

ok thanks

imqwerty (imqwerty):

np :)